OpenStudy (anikate):

A line contains the points (-4, -5), parallel to the line 5x-y=12

Parth (parthkohli):

Point slope form yo. (y)

Parth (parthkohli):

Okay, so can you convert this equation to slope-intercept form?

OpenStudy (anikate):

yup

OpenStudy (anikate):

y=5x-12

OpenStudy (anonymous):

now if a line is parllel to other their slopes are same or different????? @Anikate

OpenStudy (anikate):

same

Parth (parthkohli):

@Anikate What is the slope of a line parallel to another.

Parth (parthkohli):

Yes, and here it's 5 right? So the slope in your parallel line is also 5.

OpenStudy (anikate):

in this equation: m=5

Parth (parthkohli):

Yup. We have a point and a slope, so we'd just use the point slope form!

OpenStudy (anikate):

yes

Parth (parthkohli):

Do you know what the point-slope form is?

OpenStudy (anikate):

y=mx+b

OpenStudy (anonymous):

so a line parllel to other just have change in the value of c in y=mx+c

OpenStudy (anikate):

yes

Parth (parthkohli):

@Anikate Okay, so I'd tell you what the point-slope form of an equation is.\[\rm y - y_1 = m(x - x_1)\] if the point you are given is (x1,y1) and the slope is m.

OpenStudy (anonymous):

just make such equation with random C:)

Parth (parthkohli):

@Aperogalics No! We are also given the points!

Parth (parthkohli):

We can't just choose a random \(\rm c\).

OpenStudy (anonymous):

@ParthKohli let him make it step by step:)

OpenStudy (anonymous):

here random C doesnt mean any value..k

OpenStudy (anikate):

k i got the rest of the question thx for the help

Parth (parthkohli):

I know, @Aperogalics, C is the y-intercept. @Anikate yw!

Parth (parthkohli):

And you can't have a random y-intercept...

OpenStudy (anikate):

can u plz hepl me on my next question i will post it

OpenStudy (anonymous):

can u prove it @ParthKohli

Parth (parthkohli):

@Aperogalics A random line does NOT pass through (-4,-5).

OpenStudy (anikate):

guys i got it . can u plz help me with my next qestion

Parth (parthkohli):

Sure @Anikate!

OpenStudy (anonymous):

my frind i say c isn't mean a particular constant value i m trying to say him to write family of parllel line( general form)..:)